order nafcilin 40 mg/kg in four divided doses for sinusitis

on hand : nafcillin 250 mg scored tablets

client weight: 60 lbs

how many tablets you give per day
Answer #1

ANSWER.

* Calculation of number of tablets required per day.

     Ordered dose = 40 mg/kg (4 divided doses).

        Weight of the patient = 60 lbs.

       Convert lbs to kg for calculation.

             1 lbs = 0.454 kg.

        So

            Weight of the patient = 60 * 0.454

                                           = 27.2 kg.

            Dose needed for this patient = 40 * 27.2

                                                  = 1088 mg / day.

           Available dose = 250 mg/tablet.

Number of tablets required = 1088 / 250

                                              = 4.35

                                            = 4 tablets and 0.35 mg of one tablet.